40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 25 10:48:13 2008, Koji, Update, General, Abs. Len. Meas. ~ Realignment / beating / PLL trial 
    Reply  Mon Jul 28 15:54:04 2008, Koji, Update, General, Abs. Len. Meas. ~ More on the beat / the PLL setup PLL_setup.pngbeat_at_8MHz.jpg
       Reply  Thu Jul 31 11:19:30 2008, Koji, Update, General, Abs. Len. Meas. ~ Resonance search trial 
          Reply  Sat Aug 2 12:53:43 2008, Koji, Update, General, Abs. Len. Meas. ~ New PLL at the PSL table yarm_dc.png
             Reply  Mon Aug 4 00:37:58 2008, Koji, Update, General, Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan YEND_LAYOUT.png4th_FSR1.png4th_FSR2.png
                Reply  Mon Aug 4 05:23:57 2008, Koji, Update, General, Abs. Len. Meas. ~ Measurement for Y-arm completed 
                Reply  Mon Aug 4 21:48:24 2008, Koji, Update, General, Abs. Len. Meas. ~ Scan for TEM00/01/10 TEM00.pngTEM01.pngknife_edge.png
                   Reply  Mon Aug 4 22:31:10 2008, Koji, Update, General, Abs. Len. Meas. ~ Simple Test for TEM01/10 split TEM01_10.png
                   Reply  Tue Aug 5 00:05:57 2008, Koji, Update, General, Abs. Len. Meas. ~ IFR2023A calibration 
                      Reply  Wed Aug 6 11:10:34 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM00 scan TEM00fit.pngTEM00FSRfit.png
                         Reply  Wed Aug 6 11:43:52 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM01 scan TEM01fit.png
                            Reply  Thu Aug 14 21:52:51 2008, Koji, Summary, General, Abs. Len. Meas. ~ summary of my Summer mode_spacing_measurement_080816_v2.pdf
Message ID: 802     Entry time: Wed Aug 6 11:43:52 2008     In reply to: 801     Reply to this: 838
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ analysis of the TEM01 scan 
Analysis of the data on August 3th ~ Part 2

o I already have reported that the resonant freq of TEM10 and TEM01 split.

o Again, note that TEM10/01 were arranged almost in the horizontal/vertical by the observation of the video.

o The peaks of TEM10 and TEM01 were fitted with the same method as of TEM00.

o The peak freqs were:
f_TEM10: 5087040 Hz +/- 20 Hz
f_TEM01: 5068322 Hz +/- 15 Hz
The split is 18.7kHz.

o The additional parameter from the previous entry:
f_TEM00: 3879252 Hz +/- 9 Hz
L_yarm: 38.6462 m +/- 0.0003 m

o Radius of curvature
Rx = L /(1-Cos^2(Pi (f_TEM10 - f_TEM00) / (c/L/2) ))
Ry = L /(1-Cos^2(Pi (f_TEM01 - f_TEM00) / (c/L/2) ))

from these formula we get the value
Rx = 56.1620 +/- 0.0013 [m]
Ry = 57.3395 +/- 0.0011 [m]
Attachment 1: TEM01fit.png  58 kB  | Hide | Hide all
TEM01fit.png
ELOG V3.1.3-